OpenStudy (anonymous):

Aimee and b solved the system below 2x-3y=8 -2x+3y=-8 Aimee says the solution has no answer but b says their are infinite solutions. Who is correct? Justify

OpenStudy (jdoe0001):

look at the two equations 2x-3y=8 -2x+3y=-8 notice something the one below, is really just the one above multiplying BOTH sides by -1